### v1

# CCFExp: Facial Image Synthesis with Cycle Cross-Fusion Diffusion Model for Facial Paralysis Individuals

Abstract
Facial paralysis is a debilitating condition that affects the movement of facial muscles, leading to a significant loss of facial expressions. Currently, the diagnosis of facial paralysis remains a challenging task, often relying heavily on the subjective judgment and experience of clinicians, which can introduce variability and uncertainty in the assessment process. One promising application in real-life situations is the automatic estimation of facial paralysis. However, the scarcity of facial paralysis datasets limits the development of robust machine learning models for automated diagnosis and therapeutic interventions. To this end, this study aims to synthesize a high-quality facial paralysis dataset to address this gap, enabling more accurate and efficient algorithm training. Specifically, a novel Cycle Cross-Fusion Expression Generative Model (CCFExp) based on the diffusion model is proposed to combine different features of facial information and enhance the visual details of facial appearance and texture in facial regions, thus creating synthetic facial images that accurately represent various degrees and types of facial paralysis. We have qualitatively and quantitatively evaluated the proposed method on the commonly used public clinical datasets of facial paralysis to demonstrate its effectiveness. Experimental results indicate that the proposed method surpasses state-of-the-art methods, generating more realistic facial images and maintaining identity consistency. Code Link
22footnotetext: Corresponding author
1 Introduction
Facial palsy, which can result in temporary or permanent weakness or paralysis of one side of the face, typically presents as an acute, unilateral facial nerve weakness or paralysis that develops rapidly, often within 72 hours, with an unknown cause, affecting approximately 23 out of every 100,000 people annually [15]. Beyond the physical discomfort, patients with facial palsy often experience significant psychological distress [21]. Accurate and timely diagnosis is crucial for effective treatment; however, the variability in clinical presentation poses significant challenges. The severity estimation of facial palsy remains a largely subjective process [15].
Figure 1: Facial palsy synthesis results. Each row contains images from the same individual, and each column shows images with the same facial palsy expression.
In recent years, machine learning (ML) has shown promise in aiding the diagnosis of facial paralysis by analyzing facial features and expressions [62, 1, 14, 15]. However, the development of reliable ML models is hindered by the scarcity of comprehensive, annotated datasets that accurately represent the diversity and severity of facial paralysis. Obtaining high-quality facial images of facial palsy patients is a significant challenge due to several factors, including the need to protect patient privacy and the inherent sensitivity of medical data. Additionally, the rarity of the condition and the variability in its presentation make it difficult to gather a diverse and representative dataset. Existing datasets [64, 17, 24] are often limited in size, scope, and variability, restricting the generalization ability and performance of ML algorithms.
Several studies have explored the use of generative adversarial networks (GANs) [16] to generate facial paralysis face images [66, 50, 49]. Although [66] successfully utilizes GANs to simulate facial expressions by reconstructing facial images from extracted boundary information, the generated images often lack the fine details necessary for accurate medical diagnosis, especially when there is a significant disparity between the facial features of the patient and the face to be combined with the palsy features. Building on this, another study [50] introduces facial shape normalization to address the unnatural appearance of images resulting from shape differences between the patient and the model. While this method improves image realism, it still struggles with capturing complex facial expressions and fine details, limiting its effectiveness in more nuanced diagnostic scenarios. In addition, to leverages data augmentation, GANs are also used to improve the performance of convolutional neural networks (CNNs) in grading facial palsy [49]. This approach synthesizes additional face images with varying degrees of facial palsy to combat overfitting and enhance the model’s generalization capability. However, the synthesized images in this study primarily focus on only these five categories without adequately capturing the full range of facial expressions and subtle asymmetries inherent in real-world cases.
In this paper, we aim to address the critical need for a comprehensive facial paralysis dataset by synthesizing a large-scale, high-quality dataset. The generator is designed to capture a wide range of facial paralysis conditions and able to generate face images with varying degrees of facial paralysis based on a given facial identity image, providing a valuable resource for training ML models with the potential to improve diagnostic accuracy and clinical outcomes.
To this end, we propose a Cycle Cross-Fusion Expression Generative Model (CCFExp). CCFExp is a unique triple condition image generation model, conditioned on identity information of a given ID image, expression information and landmark information of a given style image. It could generate a realistic image of the face in the ID image exhibiting the facial palsy characteristics of the face in the style image. CCFExp employs landmark features and the corresponding loss to ensure that the model could fully capture the nuanced facial deformations, particularly the detailed changes in facial features such as eyes, mouth, and overall asymmetry. When the model is fed multiple features simultaneously, it often encounters challenges such as feature redundancy and conflicting information, which can lead to degraded performance [35]. The redundancy among features can cause the model to overweight certain aspects, while conflicting signals may lead to inconsistent learning and prediction. To address these issues, CCFExp utilize a novel cross-fusion feature integration method. This approach effectively harmonizes the information from multiple features, enabling the model to better leverage the complementary aspects of each feature, thereby significantly improving the overall performance and robustness of the model. Training diffusion models [19, 53] on limited datasets often leads to inefficiencies in learning and suboptimal model performance due to the scarcity of data. To address these challenges, we focused on improving the training efficiency and overall performance of diffusion models under constrained data conditions by modifying the training strategy. Our approach optimizes the learning process and proposes the cycle diffusion framework enabling the model to achieve superior results with a limited dataset.
Our approach extends beyond the conventional five severity levels, incorporating a wider range of facial expressions and detailed deformations of facial features, ensuring that the subtle yet critical nuances of facial movement and deformation are accurately represented. This results in a more robust model capable of generalizing to various real-world cases, thereby enhancing its clinical applicability. Moreover, this work lays the foundation for future research in automated facial analysis and the synthesis of medical datasets.
On the whole, the contributions of this paper are summarized as follows:
  • We propose a cycle framework for diffusion models, which provides an innovative approach to train higher-performance models with limited datasets.
  • We first combine diffusion models with landmark features to improve the quality of facial synthesis and introduce a clever feature confusion strategy that substantially improves the model’s capability to effectively learn from various features.
  • To address the shortage of medical datasets, our study pioneers the application of generative models in facial paralysis face synthesis while also proposing a new evaluation metric for conditionally generated objects, achieving novel SOTA results.
Figure 2: An overview of the forward process of CCFExp. a) The architecture of our feature extractors. We train the model using different images of the same patient, with one serving as the identity image $X_{ID}$ and the other as the facial palsy expression image $X_{0}$. The identity model extracts identity features from the identity image. The expression module (including $F_{Face}$ and $F_{ID}$) and $F_{LM}$ extract expression and landmark features from the facial palsy expression image, respectively. b) An illustration of the principle of the cross feature fusion strategy. We employ a method similar to cross-attention to facilitate information exchange among the three conditional features. After this interaction, the features are concatenated for further processing. c) A diagram of the noise predicting network. CCFExp utilizes a classic U-Net architecture combined with residual connections and cross-attention mechanisms to accurately predict the noise.
2 Related Work
Face Synthesis
Face generation has become a prominent area of research within computer vision, driven by advancements in generative models such as Generative Adversarial Networks (GANs) [16], Variational Autoencoders (VAEs) [31, 47, 34, 54], and diffusion models [19, 53]. Among these, GANs have been extensively utilized for various tasks such as face generation and augumentation [8, 22, 36, 44, 58, 65]. Various methods have been developed to ensure that the generated face retains the original identity while allowing for other modifications such as image styles or background [59, 4, 67, 5, 52, 46, 10, 2]. To attain a better performance, researchers have recently turned to diffusion models [23, 68, 57], which offer a more stable and probabilistic framework for face image generation. DCFace [30] proposes a dual-condition diffusion model that generates synthetic face datasets, producing high-quality face images with varied identities and styles. Our work investigates how the severity of facial palsy and changes in facial expressions influence the synthesis of realistic images.
Diffusion Models
Diffusion models [19, 53] have emerged as a robust alternative to GANs for a wide range of image generation tasks [11]. Unlike GANs, which rely on an adversarial framework, diffusion models gradually refine noisy inputs through a series of steps, leading to more stable and high-fidelity outputs. Several studies [20, 38, 51, 55, 26] have employed advanced sampling strategies that enhance both the efficiency and the output quality of diffusion models such as the latent space diffusion [48, 60, 13, 43, 40]. In addition, there are also studies [3, 28, 71, 27, 12, 32, 42, 56, 26, 29] focused on enhancing the models’ training efficiency by modifying the training strategies. We propose a novel cycle diffusion strategy that integrates the second loss setting designed to enhance model performance, particularly when training on limited datasets.
3 Method
In this section, we present our Cycle Cross-Fusion Expression Model (CCFExp) designed to synthesize high-quality facial paralysis images (see Fig. 2). First, we mention the necessary preliminary concepts about the diffusion models (sec 3.1). Next, we introduce the various feature extractors used in CCFExp which enhance the diffusion model’s ability to capture more facial details by accurately extracting multiple features (sec 3.2). In sec 3.3, we propose the cross feature fusion, which mitigates the model’s over-reliance on individual features and enhances its overall learning capability across all features by exchanging query vectors formed from different features. Finally, we devise the cycle training strategy that improves the training efficiency and performance of the diffusion model on limited datasets by modifying the training process and introducing a secondary loss function (sec 3.4).
3.1 Preliminary
Diffusion models [19, 53] are a class of generative models that define a forward process, where data is gradually corrupted by noise, and a reverse process, which learns to denoise and recover the original data distribution. The forward process is typically modeled by a discrete-time Markov chain [39], where at each time step $t$, the data $X_{t}$ is obtained by adding Gaussian noise to the data from the previous step $X_{t-1}$, following the equation:
$q(X_{t}\mid X_{t-1})=\mathcal{N}(X_{t};\sqrt{1-\beta_{t}}\,X_{t-1},\beta_{t}I),$ (1)
where $\beta_{t}$ is a variance schedule that controls the amount of noise added at each step. The reverse process aims to estimate the posterior distribution $p_{\theta}(X_{t-1}\mid X_{t})$, which is modeled by a neural network parameterized by $\theta$. The training objective is to minimize the variational bound on the negative log-likelihood, which often reduces to a simplified mean squared error loss between the true noise and the noise predicted by the model:
$L=\mathbb{E}_{t,X_{t},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(X_{t},t)\|^{2}\right],$ (2)
where $\epsilon$ is the true noise and $\epsilon_{\theta}(X_{t},t)$ is the model’s prediction. By iteratively applying this reverse process, the model is able to generate new data samples from pure noise, refining them step by step until the desired output is obtained. Building on this foundation, we employ a two-stage diffusion process along with a more complex loss function to enhance training efficiency and improve the quality of the generated outputs.
3.2 Feature Extraction Module
ID Extractor
To accurately capture identity information from images, we use the IR-50 [9], a typical CNN, as the identity feature extractor. We utilize both the output features $f_{ID}\in\mathbb{R}^{512}$ from the final layer and the intermediate layer features $I_{ID}\in\mathbb{R}^{512\times 7\times 7}$of the model to enhance the quality of image generation. For $I_{ID}$, we first flatten and rearrange their dimensions to obtain $F_{ID}\in\mathbb{R}^{49\times 512}$ that facilitates subsequent processing. Then, we concatenate the two sets of features and add positional embeddings. In summary, the output of the identity feature extraction module $E_{ID}$ is formulated as follows:
$E_{ID}(X_{ID})=Concat[f_{ID},Flatten(I_{ID})]+P_{emb}.$ (3)
For the output features, we process the intermediate feature part and the final layer feature part separately by injecting them into different model blocks, rather than concatenating them together. The identity feature extractor is trainable during the diffusion model training process to effectively extract identity features.
Expression Extractor
CCFExp employs the deviation framework of DLN [70] to accurately capture the expression features. This module consists of two models: a facial feature extraction model $F_{Face}$ and an identity feature extraction model $F_{ID}$, sharing the same network structure. Based on the assumption that facial features can be divided into identity features and expression features, we use two separate models to extract expression and identity features from the same facial image. By subtracting the identity features from the facial features, we obtain expression features that are invariant to identity. The subtracted features are then passed through a convolution operation for channel alignment. Specifically,
$E_{Exp}(X_{0})=Conv[F_{Face}(X_{0})-F_{ID}(X_{0})].$ (4)
Landmark Extractor
Facial landmarks refer to key points on a face such as the corners of the eyes, the tip of the nose, and the edges of the lips that are used to represent and understand facial geometry, which could enhance facial generation by imposing constraints on critical facial regions. For the first time, we integrate facial landmark features with diffusion models, coupled with a specifically designed loss function, to generate realistic and detailed facial images of patients with facial palsy. In our CCFExp, we employ the MobileFaceNet [6] as the landmark detector $D_{LM}$. We remove its output layer to use it as a landmark feature extractor $E_{LM}$, which outputs a feature $F_{LM}\in\mathbb{R}^{49\times 512}$.
3.3 Cross-Fusion Module
The attention mechanism [61] has become a fundamental component in many state-of-the-art models, particularly for tasks requiring the integration of information from different sources or across different parts of a data sequence. At its core, the attention mechanism allows a model to weigh the importance of different elements within a sequence, dynamically focusing on the most relevant parts to improve decision-making or feature extraction. The cross-attention mechanism extends this concept to operate between two distinct sequences. This facilitates the fusion of information from different sources, making it particularly useful for tasks that require integrating features from multiple modalities or input streams:
[…]
where $Q_{1}$ is from the first sequence and $K_{2}$, $V_{2}$ are from the second sequence.
We aim for the model to fully utilize the information from identity, expression, and landmarks when predicting noise. Although directly concatenating the features from these three sources and injecting them into the model is a feasible approach, it does not yield satisfactory results. The generated images either lose the identity information from $X_{ID}$ or fail to retain the facial expression details from $X_{0}$. Building on the concept of the cross-attention mechanism, we perform mutual interaction among multiple features, where the query sequence from one feature guides the extraction of information from the key-value sequence of another feature.
The features from the extractors are first mapped into the query, key, and value vector spaces through a QKV encoder. Specifically,
$Q=FW^{Q},K=FW^{K},V=FW^{V},$ (6)
where $F$ indicates features, $W^{Q},W^{K},W^{V}$ represent the model’s learnable parameter matrixes. Similar to the multi-head attention mechanism [61], we utilize multiple sets of transformation matrices to obtain multiple vector spaces, aiming to capture different interaction patterns among features.
Subsequently, the identity features, expression features, and landmark features each have their own corresponding QKV vector spaces. Next, we perform vector interactions. Specifically, the query vectors from the identity features $Q_{ID}$ guide the extraction of information from the expression features, the query vectors from the expression features $Q_{E}$ guide the extraction of information from the landmark features, and the query vectors from the landmark features $Q_{LM}$ guide the extraction of information from the identity features. Through Eq. 5, we obtain the three interacted features, which are then concatenated together and subsequently fed into the noise prediction network.
Figure 3: An illustration of the cycle training strategy of CCFExp. The black unidirectional arrows represent the forward diffusion process, the blue and green double-arrow dashed lines indicate the calculation processes of the first and second losses, respectively. The red dotted points represent the landmark keypoints and the brown rectangular blocks represent the extracted identity features.
3.4 Cycle Diffusion Strategy
Building upon the training framework of the original diffusion model, we incorporated an additional step of secondary diffusion. Algorithm 1 shows the training procedure of CCFExp. After the original diffusion process, where the input image $X_{0}$ is noised to generate a noisy image $X_{t}$, and the network predicts the noise $\epsilon_{\theta}$ to restore the original image $\hat{X_{0}}$. Based on the equation
$\hat{X_{0}}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(X_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(X_{t},t)\right),$ (7)
where $\bar{\alpha}_{t}$ is a term that accumulates the effects of the noise schedule over time, often defined as the product $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$, with $\alpha_{s}$ being a scalar controlling the amount of noise added at each time step, we further introduced a secondary diffusion step (see Fig. 3). In this step, the predicted image $\hat{X_{0}}$ is treated as the new input image, which is subjected to the same noise addition process to obtain a secondary noisy image $\hat{X_{t}}$. This secondary noisy image is then fed into the noise prediction network to generate a secondary predicted noise $\epsilon^{{}^{\prime}}_{\theta}$, which is subsequently used to restore the once-noised image.
For the first diffusion process, we utilized a weighted combination of mean square error loss (see Eq. 2), identity consistency loss (the same as the Time-step Dependent ID Loss in DCFace [30]), and landmark loss as our first loss function. The identity consistency loss $L_{ID}$ is derived from the cosine similarity between the identity features extracted from the images, while the landmark loss $L_{LM}$ is obtained by calculating the mean squared error of the detected landmark keypoints.
To align with the architecture of CCFExp, we introduced the concept of a secondary loss function to enhance the effectiveness of the model’s training, whose basic composition is consistent with that of the first loss. However, the noisy image $\hat{X_{t}}$ generated from $\hat{X_{0}}$ does not participate in the denoising process after the second diffusion, and $\hat{X_{0}}$ itself is also excluded from the calculation of the second loss. Instead, we calculate the respective loss components using the original image $X_{0}$ before the first diffusion and the restored predicted image $\hat{X_{0}}^{{}^{\prime}}$ after the second denoising. The final loss is the weighted sum of the first loss and the second loss.
Algorithm 1 Training Algorithm of CCFExp
[…]
for every batch do
     for $(x_{0},x_{id})$ in the batch do
         Sample $\epsilon\sim\mathcal{N}(0,I)$
         Sample $t\sim\text{Uniform}(\{0,1,2,\dots,T\})$
         Perform initial diffusion
         $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$
         Calculate $\mathcal{L}_{mse}=\|\epsilon-\epsilon_{\theta}(x_{t},x_{id},t)\|^{2}$
         $\hat{x}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}(x_{t},x_{id},t)\right)$
         Calculate $\mathcal{L}_{lm}=\|D(\hat{x}_{0})-D(x_{0})\|^{2}$
         Calculate initial identity consistency loss
         $\mathcal{L}_{id1}=CS(F(x_{id}),F(\hat{x}_{0}))$
         $\mathcal{L}_{id2}=CS(F(x_{0}),F(\hat{x}_{0}))$
         $\mathcal{L}_{id}=-\frac{t}{T}\mathcal{L}_{id1}-\frac{T-t}{T}\mathcal{L}_{id2}$
         Calculate $\mathcal{L}_{first}=\mathcal{L}_{mse}+\lambda\mathcal{L}_{id}+\gamma\mathcal{L}_{lm}$
         Perform second diffusion
         $x_{t}^{{}^{\prime}}=\sqrt{\bar{\alpha}_{t}}\hat{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$
         Calculate $\mathcal{L}^{{}^{\prime}}_{mse}=\|\epsilon-\epsilon_{\theta}^{{}^{\prime}}(x_{t}^{{}^{\prime}},x_{id},t)\|^{2}$
         $\hat{x}^{{}^{\prime}}_{0}=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}^{{}^{\prime}}(x_{t}^{{}^{\prime}},x_{id},t)\right)$
         Calculate $\mathcal{L}^{{}^{\prime}}_{lm}=\|D(\hat{x}^{{}^{\prime}}_{0})-D(x_{0})\|^{2}$
         Calculate second identity consistency loss
         $\mathcal{L}^{{}^{\prime}}_{id1}=CS(F(x_{id}),F(\hat{x}^{{}^{\prime}}_{0}))$
         $\mathcal{L}^{{}^{\prime}}_{id2}=CS(F(x_{0}),F(\hat{x}^{{}^{\prime}}_{0}))$
         $\mathcal{L}^{{}^{\prime}}_{id}=-\frac{t}{T}\mathcal{L}^{{}^{\prime}}_{id1}-\frac{T-t}{T}\mathcal{L}^{{}^{\prime}}_{id2}$
         Calculate $\mathcal{L}_{second}=\mathcal{L}^{{}^{\prime}}_{mse}+\lambda\mathcal{L}^{{}^{\prime}}_{id}+\gamma\mathcal{L}^{{}^{\prime}}_{lm}$
         Update with $\mathcal{L}=\mathcal{L}_{first}+\sigma\mathcal{L}_{second}$
     end for
end for
4 Experiments
4.1 Datasets
AFLFP:
The Annotated Facial Landmarks for Facial Palsy (AFLFP) [64] dataset is a comprehensive and laboratory-controlled collection designed to support the analysis of facial paralysis. It consists of facial images from 88 subjects, each participating in 16 different expression videos, such as brow raise, closed smile, gentle eye closure, and open smile. For each expression, the dataset captures keyframes representing four critical states: neutral, onset, a mid-state between onset and peak, and peak. In total, the AFLFP dataset contains 5,632 facial images, with 1,408 images corresponding to each key state and 64 images per subject. Each image is meticulously annotated with 68 facial landmarks, providing detailed geometric information crucial for modeling facial movements and asymmetries associated with facial paralysis.
MEEI:
The Massachusetts Eye and Ear Infirmary (MEEI) [17] dataset is a carefully curated collection focused on facial palsy, comprising both videos and high-resolution images. The dataset includes 60 videos from a diverse group of participants, consisting of 9 healthy subjects and 51 patients with varying degrees of facial paralysis. Each video captures 8 distinct facial movements, which are also provided as individual images, resulting in a total of 480 high-resolution images. The subjects in the MEEI dataset are categorized based on their eFACE scores, which assess the severity of facial palsy. The categories include normal (96–100), near-normal (91–95), mild (80–90), moderate (70–79), severe (60–69), and complete flaccid or nonflaccid facial palsy (below 60). This classification enables detailed analysis of facial movement across a broad spectrum of facial palsy severity.
Figure 4: Ablation results for the Cycle Diffusion Training Strategy, tracking training steps to calculate aFID on the test set.
4.2 Evaluation Metrics
Fréchet Inception Distance (FID) [18] is widely used to evaluate the quality of image generation. It measures the distance between the distribution of generated images and real images by comparing the statistics of feature representations extracted from a pre-trained Inception network. A lower FID score indicates that the generated images are more similar to the real images, both in terms of visual quality and diversity. However, some studies [7, 45] have shown that inconsistencies in the class distribution between the generated dataset and the real dataset can affect the calculation of the FID score. To address this issue, we propose the aligned FID (aFID) to evaluate the quality of the multi-condition diffusion model. Specifically, for each facial palsy patient in the test set, we first use images of other individuals as identity images and the patient’s condition image as the facial palsy expression image. This process synthesizes an image where the patient’s facial palsy expression is transferred onto another individual. We then use this generated dataset as the facial palsy expression images and the original facial palsy patient’s images as identity images to generate a second set of synthetic images. Finally, we calculate the FID score between this generated dataset and the real dataset. Ideally, the images in these two datasets should be identical.
Figure 5: Visualization of ablation results. CCFExp is the standard version, CCFExp1 is the CCFExp without facial landmarks, CCFExp2 excludes the cycle training strategy and CCFExp3 operates in the absence of the cross-fusion module.
4.3 Implementation details
Table 1: Quantitative results on the test set.
Method Metrics
Model GS LM CF CD aFID $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PL $\downarrow$
XID XStyle XID XStyle XID XStyle XID XStyle
DcFace 3x3 - - - 56.77 15.95 17.07 0.32 0.35 0.37 0.34 0.27 0.24
5x5 - - - 56.20 17.15 19.39 0.40 0.48 0.34 0.26 0.26 0.19
CCFExp - $\times$ $\checkmark$ $\checkmark$ 50.47 17.31 16.05 0.45 0.40 0.31 0.33 0.25 0.28
- $\checkmark$ $\times$ $\checkmark$ 54.42 17.13 15.99 0.47 0.44 0.31 0.33 0.26 0.26
- $\checkmark$ $\checkmark$ $\times$ 45.69 17.41 16.31 0.47 0.43 0.32 0.34 0.25 0.28
- $\checkmark$ $\checkmark$ $\checkmark$ 43.76 17.44 16.36 0.47 0.45 0.31 0.33 0.25 0.27
GS = Grid Size, LM = Facial Landmark Features, CF = Cross-Fusion Module, CD = Cycle Diffuion Strategy.
For the identity feature extractor $E_{ID}$, we utilized the same IR-50 [9] as DCFace [30]. Intermediate features from this network were injected into the cross-fusion module, and the final features were incorporated into the model prediction network with timesteps, and after initialization from DCFace, the model is trained end-to-end. For both the identity model and the face model within the expression extracting module, we also used the IR-50 architecture. After the same initialization, the ID model’s weights were fixed, and the module was trained on the AffectNet [41] dataset, after which the module’s weights were all fixed. For the landmark extractor $E_{LM}$ and detector $D_{LM}$, we used the MobileFaceNet [6] and fixed its weights to obtain the landmark information. For each patient’s video in the MEEI dataset, we performed frame extraction. Keyframes representing eight different expressions for each facial palsy patient were manually selected. Each expression consists of 32 frames, capturing the full expression process from onset, development, peak, dissipation, to a neutral state. For the images in AFLFP and MEEI, we first performed face extraction and segmented them into 112x112 facial images. After removing low-quality images and those with poor segmentation quality, we ultimately obtained 91,168 facial images from 313 individuals. We used 88% of the data (80,288 images from 276 individuals) for the model’s training set, while 12% of the data (10,880 images from 37 individuals) was set aside as the test set for subsequent evaluation metric calculations. CCFExp was trained on the training set using the AdamW Optimizer [33, 37] for 10 epochs with a learning rate of 0.0001, utilizing two NVIDIA RTX 4090 GPUs.
4.4 Ablation Study
To evaluate the contribution of each component in our model, we conducted a series of ablation studies. These experiments assess the impact of removing or altering key modules on the model’s performance. The visualization results of the ablation study are shown in Fig 5. To evaluate the effectiveness of image generation from different perspectives, for the synthesized images, in addition to aFID, we also compute Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) [63], Learned Perceptual Image Patch Similarity (LPIPS) [69], Perceptual Loss (PL) [25], on unseen data, by comparing them separately with the identity image XID and the facial palsy expression image XStyle. The quantitative results are shown in Tab. 1.
Facial Landmarks
To investigate the role of facial landmark features in image generation, we remove the facial landmark condition from CCFExp. The model no longer learns landmark features, and in the fusion module, only identity features and expression features are used for information exchange through query key interactions and concatenation. During training, the model also stops using the landmark loss for weight updates. The results in Tab. 1 show a significant increase in aFID, and the pixel-level comparison metrics for facial palsy expression also exhibit a clear decline. As seen in Fig. 5, the model is largely unable to correctly learn facial palsy expressions. This highlights the importance of facial landmark features in synthesizing detailed facial palsy expressions.
Cross-Fusion Module
Tab. 1 compares the performance of the model with and without the cross-fusion module. It can be observed that directly concatenating multiple features and injecting them into the model leads to a sharp increase in aFID by 10.66, making it the component with the most significant impact on overall performance. Combined with other metrics, this shows that the cross-fusion module plays a crucial role in improving both identity preservation and facial palsy feature transfer. This feature fusion module effectively enhances the model’s ability to learn from multiple features comprehensively.
Cycle Training Strategy
The design of Cycle Diffusion is intended to improve the training efficiency of the model on limited datasets. We conducted comparative experiments to verify the effectiveness of this strategy. As shown in Tab. 1, the cycle training strategy provides a noticeable performance improvement for the model. We tracked the training process and plotted the changes in aFID, with the results shown in Fig. 4. Compared to the standard diffusion model training strategy, Cycle Diffusion achieves higher training efficiency.
4.5 Comparison with Previous Methods
We compared CCFExp with a similar facial synthesis task, DCFace [30], using 3x3 and 5x5 grid sizes on unseen images. The comparison results are shown in the upper part of Tab. 1. Although DCFace 5x5 achieves better scores in several generated data and facial palsy expression data metrics, this does not necessarily indicate a higher degree of facial palsy feature transfer. Pixel-level and perception-based metrics not only focus on facial expressions but also consider factors like lighting and background, which are not primary considerations in the task of facial palsy synthesis. In terms of aFID and identity image comparison metrics, CCFExp significantly outperforms DCFace, reaching the state-of-the-art performance.
5 Conclusion
Our research proposes Cycle Cross-Fusion Expression Generative Model (CCFExp) , based on the conditional diffusion model, utilizing multiple feature extractors to accurately capture various facial features, employing a cross-fusion module to integrate these features and applying a cycle training strategy to improve training efficiency and effectiveness. As a result, CCFExp generates realistic facial palsy expression images, providing valuable training data for machine learning systems used in facial palsy diagnosis and treatment assistance. Results in Tab. 1 indicates that CCFExp surpasses existing methods in several metrics, including aFID, PSNR, and SSIM, achieving excellent performance.
Limitations
Although CCFExp has shown promising results in the field of facial palsy feature transfer, a large amount of facial palsy expression data, especially manually annotated datasets, is still needed to further enhance the model’s performance and generalization capability. We encourage more researchers to participate in the collection of facial palsy data, which will help accelerate the early application of large-scale deep learning models in this field.